OpenStudy (amonoconnor):

Would someone be willing to answer a conceptual question? My query is... Does "(sin(2x))/(cos(2x))" = "tan(2x)" ? Any and all help is greatly appreciated! :)

hartnn (hartnn):

yes thats correct

OpenStudy (amonoconnor):

Alright, then would ((sin(2x))/(cos^2(2x)) = tan^2(2x) ?

hartnn (hartnn):

no

hartnn (hartnn):

((sin^2 (2x))/(cos^2 (2x)) = tan^2 (2x)

OpenStudy (amonoconnor):

Okay... I guess my real question now is how do I got about solving this integral: ((Sin(2x))/((cos^2(2x))dx, from x=0 to pi/6 ?

OpenStudy (amonoconnor):

Can Rewrite it as: ((sin(2x))((cos(2x))^-2 ?

hartnn (hartnn):

plug in u = cos 2x :)

hartnn (hartnn):

just because you see the derivative of cos in the numerator!

OpenStudy (amonoconnor):

Right... Okay. I'm doing it! :)

hartnn (hartnn):

don't forget to change the limits too!

OpenStudy (amonoconnor):

du = -sin(2x) dx, so the integral becomes - (u^(-2))du ?

hartnn (hartnn):

du is incorrect...know chain rule ?